Darren starts a Construction business and completes the following transactions during the month of November. Create T-accounts and record transactions. Cash, A/R, Equipment, Pre-paid Rent, Un-earned Revenue, A/P, W/D, Capital, Revenue, Rent & Utilities Exp)

—a) Darren invests $40,000 of personal cash in the business

—b) He rents a small location and pays the 3 months rent of $3000.

—c) He buys $2600 of equipment and pays $1600 cash and $1000 on account

—d) He completes a landscaping job for which he is paid $1800 cash and $1200 on account.

—e) He makes a $400 payment on the money He owes.

—f) He receives a $1000 advance payment for a job he will begin soon.

—g) He has a $100 Electric bill which he hasn’t paid yet.

h) He completes the job for which he was paid in advance.

create the T accounts and record the transactions, then figure out the balance of assets, liabilities, and owner's equity.
Answer #1

T-account:

A/R 1,200 Cash 40,000 b 1,800c 1,000 e 37,800 3,000 1,600 400 Bal. Pre-paid rent 3,000 c Equipment 2,600 1 A/P Un-earned reve

Ending balances of asset, liabilities, and owners' equity are:

Assets=37,800+1,200+2,600+3,000 Liabilities =0+700 Owners equity = 40,000+4,000-100 44,600 700 43,900